Could you please help me solve the following:
4(-3/5) - 3(1/4) divided by [2(-3/5]
I came up with the answer of -213/120, but the textbook gave a different answer
(sorry, can't find the divided by symbol on the keyboard)

Found 2 solutions by akmaths, vasu2qute:
Answer by akmaths(3)

Q. is 4(-3/5) - 3(1/4 ) divided by [2(-3/5)]
soln:
First open the big bracket therefore it becomes
4(-3/5) - 3(1/4) divided by (-6/5)
Now remove the small brackets, i.e.:
-12/5 - 3/4 devided by -6/5
You can write the above as given below just by pressing
underscore key for showing the divn. line or
-(12/5 + 3/4)
_____________
- 6/5

cancel the (-) sign on top and bottom
and by finding LCM of top terms we group them together, thus
48 + 15 = 63/20 divided by 6/5
Now use multiply instead of division , that means the terms
will change places on top and bottom of '6/5' i.e
63/20 multiply by 5/6 after cancellation we get
63/24 which is the ans









Answer by vasu2qute(17)

let us solve the numerator first:
Numerator =4(-3/5)-3(1/4)= -12/5 -3/4 =(-12x4-3x5)/20 = (-48-15)/20= -63/20
Denominator =[2(-3/5)]= -6/5
Numerator/denominator =(-63/20)/(-6/5)=(-63/20)x(-5/6)=315/120=21/8